A piece of wire of cross-sectional area A and resistivity ρ is bent into a circular loop of radius r and placed in a magnetic field with its plane at right angles to the field. Determine the magnitude of the current in the loop if the field is reduced to zero in time t.
1
Expert's answer
2020-08-27T10:24:55-0400

"Given:\\A-cross-sectional\ area,\\\rho-resistivity,\\ r-radius\ of\ the\;loop,\\t-time \\Solution:\\E_i=|\frac{\Delta\Phi}{\Delta t}|=|\frac{(0-B)S}{t}|=\frac{BS}{t};\\I=\frac{E_i}{R}; R=\rho\frac{l}{A}; S=\pi r^2;\Rightarrow I=\frac{B rA}{t \rho}\\without\ taking\ into\\ account\ the\ direction\ of\ the\ current\\Answer:\\ I=\frac{B rA}{t \rho}"